# EvalAlign: Supervised Fine-Tuning Multimodal LLMs with Human-Aligned Data for Evaluating Text-to-Image Models

### v2

@@ 5.3 Comparison with Existing Evaluation Methods
SFT with human-aligned data outperforms vanilla MLLMs. To validate the effectiveness of the MLLM after SFT, we use vanilla LLaVA-NeXT 13B as the baseline model for comparison. As shown in Table 5 and Table 5, the results of vanilla model suggest some correlations with human-annotated data. However, the alignment of the vanilla MLLM is relatively low due to the absence of images generated by model (such as distorted bodies and hands images) and issues related to evaluation in the MLLM’s pre-training dataset. After applying SFT on the LLaVA-Next 13B model using human annotated data, the model’s predictions on various fine-grained evaluation metrics are almost align to the human-annotated data and significantly surpass the evaluation results of all MLLM models that were not fine-tuned. This experimental results confirms that our SFT training enables the MLLM to be successfully applied to the task of evaluating text-to-image models.
Comparison with state-of-the-art methods. To verify the human preference alignment of our model, especially when compared with other baseline methods, we calculate Kendall rank [19] and Pearson [12] correlation coefficient and summarize the results in Table 8. This comparison experiment encompasses images generated by 24 text-to-image models, covering an extensive range of synthesized image distribution.
@@ 6 Conclusion and Discussion
In this work, we design an economic and efficient evaluation method that offers high accuracy, strong generalization capabilities, and provides fine-grained, interpretable metrics. We develop a comprehensive data annotation and cleaning process tailored for evaluation tasks, and established the EvalAlign benchmark for training and evaluating models on supervised fine-tuning tasks for MLLMs. Experimental results across 24 text-to-image models demonstrate that our evaluation metrics surpass the accuracy of all the state-of-art evaluation method. Additionally, we conduct a detailed empirical study on how MLLMs can be applied to model evaluation tasks. There are still many opportunities for further advancements and expansions based on our EvalAlign. We hope that our work can inspire and facilitate future research in this field.
@@ Appendix A Limitations
Appendix B Ethical and Social Impacts
Similar with other MLLMs, EvalAlign may potentially generate responses contain offensive, inappropriate, or harmful content. Since the base MLLMs of EvalAlign are pretrained on large datasets scraped from the web that might contain private information and harmful content, they may inadvertently generate or expose sensitive information, raising ethical and privacy concerns. MLLMs are also susceptible to adversarial attacks, where inputs are intentionally crafted to deceive the model. This vulnerability can be exploited to manipulate model outputs, posing security and ethic risks. To alleviate these safety limitation, we create dedicated evaluation sets for bias detection and mitigation, and conducted adversarial testing through hours of redteaming. Besides, EvalAlign is designed for fine-grained, human-aligned automatic text-to-image evaluations. We believe that with appropriate use, it could provide users with interesting experiences for detailed synthesized image evaluation, and inspires more appealing research works about text-to-image generation.
Appendix C Annotation Details
@@
Appendix D Additional Details of the Evaluated Models
@@
Stable Diffusion {v1.4, v1.5, v2 base, v2.0, v2.1}. Stable Diffusion (SD) is a series of 1B text-to-image generative models based on latent diffusion model [41] and is trained on LAION-5B [45]. Specifically, the SD series includes SD v1.1, SD v1.2, SD v1.4, SD v1.5, SD v2 base, SD v2.0, and SD v2.1 respectively. Among them, we choose the most commonly-employed SD v1.4, SD v1.5, SD v2.0 and SD v2.1 for EvalAlign evaluation.
@@
Latent Consistency Model Stable Diffusion XL Latent Consistency Model Stable Diffusion XL (LCM SDXL) [28] enables SDXL for swift inference with minimal steps. Viewing the guided reverse diffusion process as solving an augmented probability flow ODE (PF-ODE), LCMs are designed to directly predict the solution of such ODE in latent space, mitigating the need for numerous iterations and allowing rapid, high-fidelity sampling.
Dreamlike Diffusion 1.0. Dreamlike Diffusion 1.0 [10] is a SD v1.5 model finetuned on high-quality art images by dreamlike.art.
@@
Dreamlike Photoreal 2.0. Dreamlike Photoreal 2.0 [11] is a photorealistic text-to-image latent diffusion model resumed from SD v1.5 by dreamlike art. This model was finetuned on 768x768 images, it works pretty good with resolution 768x768, 640x896, 896x640 and higher resolution such as 768x1024.
@@
Redshift Diffusion. Redshift Diffusion [39] is a Stable Diffusion model finetuned on high-resolution 3D artworks.
@@
Vintedois Diffusion. Vintedois Diffusion [51] is a Stable Diffusion v1.5 model finetuned on a large number of high-quality images with simple prompts to generate beautiful images without a lot of prompt engineering.
@@
Safe Stable Diffusion {Weak, Medium, Strong, Max}. Safe Stable Diffusion [31] is an enhanced version of the SD v1.5 model by mitigating inappropriate degeneration caused by pretraining on unfiltered web-crawled datasets. For instance SD may unexpectedly generate nudity, violence, images depicting self-harm, and otherwise offensive content. Safe Stable Diffusion is an extension of Stable Diffusion that drastically reduces this type of content. Specifically, it has an additional safety guidance mechanism that aims to suppress and remove inappropriate content (hate, harassment, violence, self-harm, sexual content, shocking images, and illegal activity) during image generation. The strength levels for inappropriate content removal are categorized as: {Weak, Medium, Strong, Max}.
@@
MultiFusion. MultiFusion [29] is a multimodal, multilingual diffusion model that extends the capabilities of SD v1.4 by integrating various modules to transfer capabilities to the downstream model. This combination results in novel decoder embeddings, which enable prompting of the image generation model with interleaved multimodal, multilingual inputs, despite being trained solely on monomodal data in a single language.
@@
DeepFloyd-IF { M, L, XL } v1.0. DeepFloyd-IF [1] is a novel state-of-the-art open-source text-to-image model with a high degree of photorealism and language understanding. It is a modular composed of a frozen text encoder and three cascaded pixel diffusion modules: a base model that generates 64x64 image based on text prompt and two super-resolution models, each designed to generate images of increasing resolution: 256x256 and 1024x1024, respectively. All stages of the model utilize a frozen text encoder based on the T5 transformer to extract text embeddings, which are then fed into a UNet architecture enhanced with cross-attention and attention pooling. Besides, it underscores the potential of larger UNet architectures in the first stage of cascaded diffusion models and depicts a promising future for text-to-image synthesis. The model is available in three different sizes: M, L, and XL. M has 0.4B parameters, L has 0.9B parameters, and XL has 4.3B parameters.
Throughout all of the conducted experiments, we use the default inference settings for each evaluated models.
Appendix E Instruction Templates
@@
Appendix F Additional Quantitative Analysis
F.1 Generalization Experiments
@@
F.2 Instruction Enhancement Experiments
@@
F.3 Mulit-scaling Resolutions Experiments
@@
F.4 Full Comparison with Existing Evaluation Methods
Model | Human | EvalAlign | HPS v2 | CLIP-score | ImageReward | PickScore | IS |
PixArt XL2 1024 MS [6] | 2.2848 | 1.6415 | 31.6226 | 0.8580 | 0.9696 | 22.1335 | 5.258321 |
Dreamlike Photoreal v2.0 [11] | 2.0070 | 1.4522 | 29.2322 | 0.828612 | 0.1886 13 | 21.2271 | 5.645214 |
SDXL Refiner v1.0 [47] | 1.9229 | 1.6072 | 29.8197 | 0.8566 | 0.7245 | 22.0492 | 5.398516 |
SDXL v1.0 [33] | 1.8136 | 1.4675 | 29.0620 | 0.8467 | 0.7043 | 21.8106 | 5.377418 |
Wuerstchen [32] | 1.7837 | 1.4279 | 30.6622 | 0.819914 | 0.3212 11 | 21.3720 | 5.9751 |
LCM SDXL [28] | 1.6910 | 1.3391 | 29.3588 | 0.833510 | 0.5304 | 21.6532 | 5.9874 |
Openjourney [34] | 1.6667 | 1.175010 | 26.347513 | 0.819615 | 0.1478 16 | 20.863710 | 5.736810 |
Safe SD MAX [31] | 1.6491 | 1.2175 | 25.739617 | 0.755524 | -0.050722 | 20.459421 | 5.542815 |
LCM LORA SDXL [28] | 1.6387 | 1.3833 | 27.329910 | 0.8364 | 0.4959 | 21.4824 | 5.657512 |
Safe SD STRONG [31] | 1.630810 | 1.146611 | 25.576418 | 0.816518 | -0.102223 | 20.621118 | 5.8643 |
Safe SD MEDIUM [31] | 1.627511 | 1.129815 | 26.279814 | 0.810120 | 0.2042 12 | 20.788012 | 6.1760 |
Safe SD WEAK [31] | 1.607812 | 1.118817 | 26.118015 | 0.780923 | -0.126424 | 20.387324 | 5.386117 |
SD v2.1 [40] | 1.552413 | 1.109418 | 26.582312 | 0.8377 | 0.4116 | 21.0502 | 5.307319 |
SD v2.0 [40] | 1.527714 | 1.130014 | 25.348121 | 0.817017 | 0.0872 18 | 20.752913 | 5.9060 |
Openjourney v2 [35] | 1.500015 | 0.995620 | 24.698423 | 0.795822 | -0.041521 | 20.408822 | 5.199522 |
Redshift diffusion [39] | 1.473316 | 1.138212 | 25.157222 | 0.810121 | 0.0218 20 | 20.615519 | 5.7657 |
Dreamlike Diffusion v1.0 [10] | 1.465217 | 1.2052 | 29.6506 | 0.8543 | 0.6508 | 21.2664 | 5.661411 |
SD v1.5 [40] | 1.441718 | 1.136213 | 25.497219 | 0.821413 | 0.1686 14 | 20.714316 | 6.0535 |
IF-I-XL v1.0 [1] | 1.380819 | 0.922122 | 27.4512 | 0.8449 | 0.6087 | 20.747414 | 5.301220 |
SD v1.4 [40] | 1.359220 | 0.951121 | 25.369720 | 0.819016 | 0.1050 17 | 20.653517 | 5.8571 |
Vintedois Diffusion v0.1 [51] | 1.356221 | 1.079719 | 26.590111 | 0.8341 | 0.3562 10 | 20.835811 | 5.018123 |
IF-I-L v1.0 [1] | 1.263522 | 0.881423 | 27.4836 | 0.8384 | 0.4463 | 20.717015 | 5.650213 |
MultiFusion [29] | 1.237223 | 1.129816 | 23.813324 | 0.815119 | 0.0695 19 | 20.478020 | 4.382424 |
IF-I-M v1.0 [1] | 1.013524 | 0.792824 | 25.952216 | 0.832911 | 0.1637 15 | 20.403523 | 5.7451 |
Model | Human | EvalAlign | HPS v2 | CLIP-score | ImageReward | PickScore | IS |
IF-I-XL v1.0 [1] | 5.4500 | 5.5300 | 32.547710 | 0.8579 | 0.4391 | 21.199810 | 7.1864 |
IF-I-L v1.0 [1] | 5.2300 | 5.4500 | 32.7140 | 0.8538 | 0.3820 | 21.128412 | 6.657120 |
SDXL Refiner v1.0 [47] | 5.2100 | 5.4000 | 35.6465 | 0.8528 | 0.4738 | 22.3532 | 7.1101 |
LCM SDXL [28] | 5.1800 | 5.3300 | 33.8011 | 0.8512 | 0.3833 | 21.9620 | 7.006714 |
PixArt XL2 1024 MS [6] | 5.1100 | 5.3100 | 37.0493 | 0.8634 | 0.6542 | 22.3926 | 6.929316 |
IF-I-M v1.0 [1] | 5.0800 | 5.2200 | 31.095114 | 0.8434 | 0.0499 10 | 20.827020 | 7.2586 |
LCM LORA SDXL [28] | 5.0600 | 5.2700 | 32.7752 | 0.834910 | 0.1618 | 21.7627 | 6.838919 |
SDXL v1.0 [33] | 5.0300 | 5.3500 | 35.1593 | 0.8540 | 0.4322 | 22.1291 | 7.2762 |
Wuerstchen [32] | 4.8700 | 5.1700 | 36.4632 | 0.8381 | 0.2513 | 21.7779 | 7.1280 |
Openjourney [34] | 4.830010 | 4.920015 | 31.149512 | 0.817316 | -0.086714 | 21.116313 | 6.272922 |
SD v2.1 [40] | 4.800011 | 5.070011 | 31.101713 | 0.827814 | -0.045312 | 21.2093 | 7.0767 |
MultiFusion [29] | 4.680012 | 4.800018 | 28.795724 | 0.826415 | -0.133715 | 20.962517 | 4.959324 |
Dreamlike Diffusion v1.0 [10] | 4.660013 | 5.150010 | 34.8196 | 0.8493 | 0.2295 | 21.5550 | 6.858118 |
SD v2.0 [40] | 4.640014 | 5.010012 | 30.615317 | 0.829813 | -0.142416 | 21.190511 | 7.012413 |
Vintedois Diffusion v0.1 [51] | 4.620015 | 4.950014 | 31.950311 | 0.831912 | -0.022211 | 21.114114 | 6.198723 |
Safe SD STRONG [31] | 4.600016 | 4.830017 | 30.661516 | 0.775123 | -0.502822 | 20.749121 | 7.2743 |
Dreamlike Photoreal v2.0 [11] | 4.560017 | 4.980013 | 33.7712 | 0.834411 | -0.085913 | 21.4832 | 6.936315 |
Safe SD WEAK [31] | 4.530018 | 4.710020 | 30.564418 | 0.814018 | -0.272818 | 20.989916 | 6.423821 |
SD v1.4 [40] | 4.520019 | 4.760019 | 29.914920 | 0.804820 | -0.343819 | 20.846219 | 7.015012 |
SD v1.5 [40] | 4.450020 | 4.900016 | 30.167319 | 0.814217 | -0.221317 | 20.864018 | 7.1642 |
Safe SD MEDIUM [31] | 4.400021 | 4.560024 | 30.782015 | 0.797421 | -0.359120 | 21.025715 | 7.070910 |
Redshift diffusion [39] | 4.350022 | 4.670021 | 29.286522 | 0.806619 | -0.417221 | 20.632723 | 7.061011 |
Safe SD MAX [31] | 4.310023 | 4.590023 | 29.812621 | 0.760124 | -0.609524 | 20.704622 | 7.2273 |
Openjourney v2 [35] | 4.150024 | 4.650022 | 29.238923 | 0.785122 | -0.605123 | 20.597324 | 6.861317 |
Table 14 and Table 15 present the comparison results between our proposed method and existing alternatives. Owing to the powerful aligned understanding of images and text by MLLMs, our model achieves the best alignment with human performance across all tested benchmarks for 24 text-to-image models, along with a significant improvement in image faithfulness. This indicates that our model excels in both comprehending and evaluating the intricate details of generated images, closely mirroring human judgment and setting a new standard for image-text alignment and faithfulness.
Appendix G Qualitative Analysis
@@
As shown in Figure 5, we present a comparison of different evaluation metrics on images generated by four models, including human annotated scores, EvalAlign, ImageReward [56], HPSv2 [54], and PickScore [20]. The digits in the figure represent the ranking for each evaluation metric, with darker colors indicating higher rankings. From the figure, it is evident that our proposed EvalAlign metric closely matches the human rankings across two evaluation dimensions, demonstrating excellent consistency.
Furthermore, Figure 6 showcases the EvalAlign evaluation metric across different fine-grained results. The numbers in the figure represent EvalAlign scores for the corresponding evaluation aspect, with darker colors indicating higher scores and better generation performance. Note that if the text prompt does not specify a particular style, the style consistency score defaults to 0. From these results, it is evident that the same text-to-image model exhibits significant performance variation across different evaluation aspects.

### v3

@@ 5.3 Comparison with Existing Evaluation Methods
SFT with human-aligned data outperforms vanilla MLLMs. To validate the effectiveness of the MLLM after SFT, we use vanilla LLaVA-NeXT 13B as the baseline model for comparison. As shown in Table 5 and Table 5, the results of vanilla model suggest some correlations with human-annotated data. However, the alignment of the vanilla MLLM is relatively low due to the absence of images generated by model (such as distorted bodies and hands images) and issues related to evaluation in the MLLM’s pre-training dataset. After applying SFT on the LLaVA-Next 13B model using human annotated data, the model’s predictions on various fine-grained evaluation metrics are almost align to the human-annotated data and significantly surpass the evaluation results of all MLLM models that are not finetuned. This experimental results confirms that our SFT training enables the MLLM to be successfully applied to the task of evaluating text-to-image models.
Comparison with other methods. To verify the human preference alignment of our model, especially when compared with other baseline methods, we calculate Kendall rank (KENDALL, 1938) and Pearson (Freedman et al., 2007) correlation coefficient on images generated by 24 text-to-image models and summarize the results in Table 8.
@@ 6 Conclusion and Discussion
In this work, we design an economic evaluation method that offers high accuracy, strong generalization capabilities, and provides fine-grained, interpretable metrics. We develop a comprehensive data annotation and cleaning process tailored for evaluation tasks, and establish the EvalAlign benchmark for training and evaluating models on supervised fine-tuning tasks for MLLMs. Experimental results across 24 text-to-image models demonstrate that our evaluation metrics surpass the accuracy of all the state-of-art evaluation method. Additionally, we conduct a detailed empirical study on how MLLMs can be applied to model evaluation tasks. There are still many opportunities for further advancements and expansions based on our EvalAlign. We hope that our work can inspire and facilitate future research in this field.
7 Reproducibility Statement
The full version of the source code, dataset, as well as the final version of the finetuned MLLMs (one finetuned on LLaVA-NeXT 13B and the other one finetuned on LLaVA-NeXT 34B) will be released to the public. The data construction procuedure, including data collection and curation, data cleaning and annotation, is thoroughly described in Section 3. For details related to the human annotation and the measures that used to ensure its quality, we comprehensively introduce them in Appendix B. As for every experiment introduced in this paper, we provide a general introduction in Section 5 and exhibit implementation details related to reproduce our experiments. Specifically, the latter includes the hyper-parameters of each evaluated models, the employed instruction, as well as more supplementary experiments, which are described in Appendix C, Appendix D and Appendix E.
8 Ethics Statement
We are committed to conducting this research with the highest ethical standards. Our goal is to contribute positively to the fields of evaluation benchmarks on artificial intelligence generated content, emphasizing transparency and reproducibility in our design. Similar with other MLLMs, EvalAlign may potentially generate responses contain offensive, inappropriate, or harmful content. Since the base MLLMs of EvalAlign are pretrained on large datasets scraped from the web that might contain private information and harmful content, they may inadvertently generate or expose sensitive information, raising ethical and privacy concerns. MLLMs are also susceptible to adversarial attacks, where inputs are intentionally crafted to deceive the model. This vulnerability can be exploited to manipulate model outputs, posing security and ethic risks. To alleviate these safety limitation and our fulfill our social responsibility as artificial intelligence researchers, we create dedicated evaluation sets for bias detection and mitigation, and conducted adversarial testing through hours of redteaming. Besides, EvalAlign is designed for fine-grained, human-aligned automatic text-to-image evaluations, which can serve as a stepping stones toward revealing the inner generation nature of text-to-image generative models, thereby lowering the ethical hazard of these models. We believe that with appropriate use, it could provide users with interesting experiences for detailed synthesized image evaluation, and inspires more appealing research works about text-to-image generation.
@@ Appendix A Limitations
Appendix B Annotation Details
@@
Appendix C Additional Details of the Evaluated Models
@@
Stable Diffusion {v1.4, v1.5, v2 base, v2.0, v2.1}. Stable Diffusion (SD) is a series of 1B text-to-image generative models based on latent diffusion model (Rombach et al., 2022b) and is trained on LAION-5B (Schuhmann et al., 2022). Specifically, the SD series includes SD v1.1, SD v1.2, SD v1.4, SD v1.5, SD v2 base, SD v2.0, and SD v2.1 respectively. Among them, we choose the most commonly-employed SD v1.4, SD v1.5, SD v2.0 and SD v2.1 for EvalAlign evaluation.
@@
Latent Consistency Model Stable Diffusion XL Latent Consistency Model Stable Diffusion XL (LCM SDXL) Luo et al. (2023) enables SDXL for swift inference with minimal steps. Viewing the guided reverse diffusion process as solving an augmented probability flow ODE (PF-ODE), LCMs are designed to directly predict the solution of such ODE in latent space, mitigating the need for numerous iterations and allowing rapid, high-fidelity sampling.
Dreamlike Diffusion 1.0. Dreamlike Diffusion 1.0 (dreamlike.art, a) is a SD v1.5 model finetuned on high-quality art images by dreamlike.art.
@@
Dreamlike Photoreal 2.0. Dreamlike Photoreal 2.0 (dreamlike.art, b) is a photorealistic text-to-image latent diffusion model resumed from SD v1.5 by dreamlike art. This model was finetuned on 768x768 images, it works pretty good with resolution 768x768, 640x896, 896x640 and higher resolution such as 768x1024.
Openjourney v1, v2. Openjourney (PromptHero, a) is an open-source text-to-image generation model resumed from SD v1.5 and finetuned on Midjourney images by PromptHero. Openjourney v2 (PromptHero, b) was further finetuned using another 124000 images for 12400 steps, about 4 epochs and 32 training hours.
@@
Redshift Diffusion. Redshift Diffusion (Redshift-Diffusion, ) is a Stable Diffusion model finetuned on high-resolution 3D artworks.
@@
Vintedois Diffusion. Vintedois Diffusion (Vintedois-Diffusion v0., 1) is a Stable Diffusion v1.5 model finetuned on a large number of high-quality images with simple prompts to generate beautiful images without a lot of prompt engineering.
@@
Safe Stable Diffusion {Weak, Medium, Strong, Max}. Safe Stable Diffusion (Patrick et al., 2022) is an enhanced version of the SD v1.5 model by mitigating inappropriate degeneration caused by pretraining on unfiltered web-crawled datasets. For instance SD may unexpectedly generate nudity, violence, images depicting self-harm, and otherwise offensive content. Safe Stable Diffusion is an extension of Stable Diffusion that drastically reduces this type of content. Specifically, it has an additional safety guidance mechanism that aims to suppress and remove inappropriate content (hate, harassment, violence, self-harm, sexual content, shocking images, and illegal activity) during image generation. The strength levels for inappropriate content removal are categorized as: {Weak, Medium, Strong, Max}.
@@
MultiFusion. MultiFusion (Marco et al., 2023) is a multimodal, multilingual diffusion model that extends the capabilities of SD v1.4 by integrating various modules to transfer capabilities to the downstream model. This combination results in novel decoder embeddings, which enable prompting of the image generation model with interleaved multimodal, multilingual inputs, despite being trained solely on monomodal data in a single language.
@@
DeepFloyd-IF { M, L, XL } v1.0. DeepFloyd-IF (Alex Shonenkov & et al., 2023) is a novel state-of-the-art open-source text-to-image model with a high degree of photorealism and language understanding. It is a modular composed of a frozen text encoder and three cascaded pixel diffusion modules: a base model that generates 64x64 image based on text prompt and two super-resolution models, each designed to generate images of increasing resolution: 256x256 and 1024x1024, respectively. All stages of the model utilize a frozen text encoder based on the T5 transformer to extract text embeddings, which are then fed into a UNet architecture enhanced with cross-attention and attention pooling. Besides, it underscores the potential of larger UNet architectures in the first stage of cascaded diffusion models and depicts a promising future for text-to-image synthesis. The model is available in three different sizes: M, L, and XL. M has 0.4B parameters, L has 0.9B parameters, and XL has 4.3B parameters.
Appendix D Instruction Templates
@@
Appendix E Additional Quantitative Analysis
E.1 Generalization Experiments
@@
E.2 Instruction Enhancement Experiments
@@
E.3 Mulit-scaling Resolutions Experiments
@@
Appendix F Qualitative Analysis
@@
As shown in Figure 5, we present a comparison of different evaluation metrics on images generated by four models, including human annotated scores, EvalAlign, ImageReward (Xu et al., 2024), HPSv2 (Wu et al., 2023a), and PickScore (Kirstain et al., 2024). The digits in the figure represent the ranking for each evaluation metric, with darker colors indicating higher rankings. From the figure, it is evident that our proposed EvalAlign metric closely matches the human rankings across two evaluation dimensions, demonstrating excellent consistency. Specifically, the numbers in the figure represent EvalAlign scores for the corresponding evaluation aspect, with darker colors indicating higher scores and better generation performance. Note that if the text prompt does not specify a particular style, the style consistency score defaults to 0. From these results, it is evident that the same text-to-image model exhibits significant performance variation across different evaluation aspects.